Sol Summary – November 9th

Sol: 2

Summary Title: Rain on Mars???

Author’s name: Madison Rooth

Mission Status: Nominal

Sol Activity Summary: This morning our crew gathered around our kitchen table and discussed our EVA plans over a breakfast of coffee and Cliff bars. EVA Team 1, consisting of CO Crisman, XO Rooth, and MCA Murphy, headed out via rovers Curiosity and Perseverance to Kissing Camel Ridge on the East side at 0900 to sample the regolith for later analysis. Team 1 returned at 1200 and relaxed over a lunch of cheese and broccoli soup.

EVA Team 2 consisting of GHO Montanez, ENG Moga, and HSO Milne embarked for their EVA to the West side of Kissing Camel Ridge at 1300. However, though the wind started at 10 mph at EVA start, it increased to around 15 mph shortly after embarkation with light rain while on EVA, so Team 2 turned around per EVA Briefing and returned to Hab. They arrived back at base at 1340.

After this, the crew worked on individual research for the rest of the day.

Look Ahead Plan: Conduct EVA to West side of Kissing Camel Ridge for EVA Team 2

Anomalies in work: Poor weather lead to an aborted EVA

Weather: Windy with a little bit of rain

Crew Physical Status: All good, 1 minor abrasion during morning EVA

EVA: 2 to Kissing Camel Ridge, the second one was aborted after about 30 minutes

Reports to be filed: Operations Report, HSO report, Images for SOL 1 and SOL 2

Support Requested: None at this time

Operations Report – November 9th

Crew 239 Operations Report 9-11-2021

Name of person filing report: Aidan Moga

Non-nominal systems: toilet, generator

Action taken for non-nominal systems: Toilet continues to leak from fitting, the temporary solution of closing the valve has successfully reduced the water leakage. The generator is still down, so power is continued to be conserved. Due to the cloudy weather, the power was more rationed. As per Outpost’s request, the heating was turned off, along with all nonessential lights and unused crew suits. Additionally, all device charging was minimized, and rovers weren’t charged according to power-saving protocols.

Daily crew water use: 20 gallons

GreenHab water use: 4 gallons

Hab toilet tank emptied: Yes

Rovers used: Yes

Notes on rovers (report nominal or any anomalies): The ignition in Curiosity was loose, making it more difficult to start the rover during EVAs. Will do an engineering EVA tomorrow, in better weather.

Summary of Crew and Mission Support internet: Nominal

Summary of suits and radios: Nominal

Campus-wide inspection, if action is taken, what and why?: Nominal

Summary of general operations: The tunnel system was inspected

Summary of Hab operations: The Hab toilet tank was emptied.

Summary of GreenHab operations (note any harvested crops):

Temperature throughout Sol 2:

9:31 a.m. -> 70.4 degree F with 19% humidity
2:22 p.m. -> 71.1 degree F with 19% humidity
5:51 p.m. -> 67.3 degree F with 20% humidity
Hours of Supplemental Light:

From 6:00 p.m. – 7:00 p.m. for watering, transplanting, and cleaning the GreenHab

Time(s) of watering for crops: 6:00 p.m.


Organization and cleaning the GreenHab happened most of the day. I rearranged all the pots and placed them along with the watering cans and stakes on the front most shelf of the hab.

Harvested 0.75 lbs of rosemary

We plan on making either mozzarella rosemary bread or a rosemary garlic butter

Transplanted the mint and rosemary plants from the herb bed to the hanging pots closest to the blue tank. Since these plants are known to ‘take over’ a garden, it is best to separate these plants into their own pots for better control.

All the zucchini’s have a disease on the leaves that cannot be identified. Need assistance on how to handle this (images attached to this email).

The crew measured the blue tank to figure out how much liquid it can hold and calculated it to be ~242 gallons. We figured it out by measuring the volume of a dodecagon cylinder.

Support/ Supplied Needed: Would it be possible to have a type of manure?

Summary of ScienceDome operations: Nominal

Summary of RAM operations: Nominal

Summary of any observatory issues: NA

Summary of health and safety issues: HSO report filed

Questions, concerns, supplies needed, and requests:

The crew is requesting some waste removal. Burnable and non-burnable waste was placed in the rear airlock. Additionally, the oil change receipt was put in the airlock.

Journalist Report – November 9th

Journalist Report 09 Nov 2021
By Madison Rooth and Brian Murphy

MDRS Crew 239 Sol 2 (11/09/2021):

We began our morning with a breakfast of water, coffee, and Cliff bars while preparing for our EVA at 0900. The crew studied our Martian Map and began suiting up for EVA Team 1 to embark on their mission. CO Crisman led the EVA, and XO Rooth and MCA Murphy joined with goals of geologic sampling and rover training.

We had a very productive EVA with samplings of sediments, marble, and local regolith to analyze later. MCA Murphy is excited to analyze the grain types in the regolith for their compatibility in clay pottery.

Upon Team 1’s return around 1200, the crew shared a meal of broccoli and cheddar soup with rice.

Then, Team 2, consisting of GHO Montanez, ENG Moga, and HSO Milne, suited up to begin their EVA to the West side of Kissing Camel Ridge. At this time, however, dark clouds were appearing on the horizon, but the wind was still holding steady at 10 mph. CO Crisman informed EVA Team 2 Lead GHO Montanez that if the wind speeds faster than this, to turn around and return to Hab. About half an hour after leaving, MDRS Director Dr. Shannon Rupert contacted the Hab Support Team and informed them of the dangerous weather situation.

However, EVA Team 2 was already out of the radio contact perimeter. Luckily, according to the team briefing earlier, with the winds now speeding at around 15 mph, EVA Team 2 had already begun to return to Hab and safely arrived only at 1330.

While this is a disappointing setback, as Team 2 was excited to explore the rocky Martian terrain, we are preparing to reschedule their EVA to the following morning when the weather will (hopefully) be more cooperative with human exploration of the Red Planet. The safety of our crew must always come first.

When Team 2 returned, the crew worked on individual research projects. GHO Montanez trekked through the Hab Tunnels to the Green House to care for the plants. ENG Moga and HSO Milne worked to ensure a safe and comfortable Hab environment, while XO Rooth and MCA Murphy worked with sediment samples.

Since the generator is still misbehaving, the Hab, unfortunately, does not have power (including heat!!), but this does not discourage our astronauts from putting their best boot forward in this dusty terrain. The crew worked until dinnertime, where GHO Montanez and HSO Milne prepared a dinner of Au Gratin potatoes and meat from the food resupply probe, delivered in the evening.

All in all, Crew 239 continues to push humanity further (and farther) than it has gone before and is looking forward to the following Sol when the Martian atmosphere will not be as harsh.

Crew 239 Operations Report for November 8th

Crisman, Keith <keith.crisman>

9:48 PM (11 minutes ago)

MDRS Operations Report 8Nov2021

Name of person filing report: Aidan Moga

Reason for Report: Daily

CrewCar Used: Yes, the CrewCar was used to transport the crew from Grand Junction to MDRS. The CrewCar was disinfected. The recorded mileage was 197343. The car was fueled and inspected for damage. The window visors are broken on the driver’s side and rear passenger side. There is a missing hubcap on the driver’s front side. The oil was changed at Walmart as requested. The oil is at 1/8-1/4 below the top of safe, the left tail light is out sometimes, brake and coolant levels are within safe range. The car was refueled at hollow mountain for 10.21 gallons or $40.89.

Non-nominal systems: Toilet, Generator

Action taken for non-nominal systems: Toilet was leaking inside of the water inlet fitting. We attempted to fix the leak by removing the fitting, cleaning the threads, placing plumbers’ tape and tightening the fitting. The leak source appears to be from inside the toilet control mechanism where the pedal actuates the valve. Temporary solution is to turn off water to the toilet using the valve behind the toilet when not in use. The toilet also smells bad during the high winds. Generator is still waiting for repairs to the propane supply. Electricity will continue to be rationed after 2PM.

ScienceDome Dual Split: NA

Solar – Nominal, supplying all power

Notes on solar: None

Water (loft tank): 50 Gallons

Water (static tank) – Have two of three shipments from the HABCar transfer tank. The third shipment will come on Thursday.

Static to Loft Pump used – Yes

Water in GreenHab: Around 40% full

Water in Science Dome: 0 gallons

Hab toilet tank emptied: No

Perseverance rover used: No

Sojourner rover used: No

Spirit rover used: No

Opportunity rover used: No

Curiosity rover used: No

Notes on rovers: Rovers were unplugged to conserve power after 2pm due to generator issues.

ATV’s Used: No

Summary of internet: Nominal

Summary of suits and radios: Suits were assigned to crew members. One suit had issues with pulling a crew member’s hair at the front of the suit.

Summary of general operations: Checked the tunnel system, added a few zip ties to secure the tunnel system.

Summary of Hab operations: Adhered to power rationing protocols, set heating to 58 degrees for tonight.

Summary of Greenhab operations: Nothing to report

Summary of Sciencedome operations: Nothing to report

Summary of RAM operations: Failed to locate the RAM fire alarm that was mentioned in the HSO checkbook.

Summary of observatory issues: None

Summary of health safety issues: None

Questions, concerns, supplies needed and requests: None

Crew 239 Sol Summary for November 8th

Sol: 1

Summary Title: First Day of Martian Activity!

Author’s name: Madison Rooth

Mission Status: Nominal

Sol Activity Summary: Our crew awoke this morning and shared a breakfast of delicious pancakes cooked by Crew Artist Brian Murphy. After this, EVA Team 1 of CO Keith Crisman, XO Madison Rooth, and MCA Brian Murphy embarked to Marble Ritual for a geologic sampling and training EVA at 1000.Team 1 arrived back at the Hab at 1059 and cooled off while preparing EVA Team 2 for their mission to Marble Ritual. EVA Team 2 left Hab at 1332 and consisted of GHO Cynthia Montanez, HSO Emily Milne, and ENG Aidan Moga. Following EVA Team 2’s return to the Hab at 1500, we shared a delicious meal of rice, vegetables, and chicken. ENG Aidan and CO Crisman attempted to fix our waste facility (toilet) due to a freshwater leak). The crew then worked on individual research until dinnertime.

Look Ahead Plan: Conduct EVAs to Kissing Camel Ridge

Anomalies in work: No anomalies

Weather: Sunny and hot

Crew Physical Status: All good

EVA: 2 to Marble Ritual

Reports to be filed: Operations report on it’s way!

Support Requested: None at this time

Crew 239 GreenHab Report for November 8th

Crew 239 GreenHab Report 08-11-2021

GreenHab Officer: Cynthia Montanez

Environment Control: Opened door at 11:31 a.m. when the temperature in the GreenHab was 88℉

Average Temperatures:

1. Temperature at 8:30 a.m. à 67℉ with 8% humidity

2. Temperature at 11:31 a.m. à 88℉ with 14% humidity

3. Temperature at 11:31 a.m. à 70℉ with 19% humidity

Hours of supplemental light: None

Daily water usage for crops: One full white watering can

– Need precise measurements of pail volume for proper values

Daily water usage for research and/or other purposes: None

Water in the Blue Tank ___0___ gallons.

– Need the amount of gallons the blue tank can hold water

Time(s) of watering for crops: 5:33 p.m.

Changes to crops: None

Narrative: Plants are growing well. Needed to leave the GreenHab door open most of the day since it would not cool down until the Sun went down.


1. Fresh Radish (12.09 lbs)

2. Mixed Greens (2.45 lbs)

3. Wild Rocket (0.65 lbs)

Support/ Supplied Needed: Soil

Crew 239 Sol Summary for November 7th

Crew 239 Sol Summary Report 11-07-2021

Sol: 1

Summary Title: Beginning of sim

Author’s name: Madison Rooth

Mission Status: Nominal

Sol Activity Summary: We arrived in Hollow Mountain to fuel the CrewCar at 10:55am and arrived at the Hab at 11:45am. We conducted an overview of the site and met with Shannon shortly after. After system introductions and acclimating the Crew to their new surroundings, two members, XO Madison Rooth and ENG Aidan Moga followed Shannon into Hanksville to fill the HabCar tank with water from Hollow Mountain. Due to difficulties with the hose and pump, it took several hours to pump the water into the Static Tank. After this was completed, HSO Emily Milne and GHO Cynthia Montanez made the second trip into town to fill the HabCar tank and transfer it to the Static Tank. CO Keith Crisman, MCA Brian Murphy, and XO Madison Rooth repaired the window in the Hab’s main door due to it breaking from the high winds.

Look Ahead Plan: Conduct Marble Ritual EVAs

Anomalies in work: No anomalies

Weather: Sunny and windy

Crew Physical Status: All good


Reports to be filed: HSO Report

Support Requested: None at this time

Copyright © The Mars Society. All rights reserved. | Main Site